# Can repeating and non-repeating FRBs be drawn from the same population?

### v1

@@ 3 Varying activity rates
Crucially, the assumption of a PL rate distribution is directly supported by observations. Lu et al. (2022) estimated the number density of several empirically defined FRB sub-classes, based on the distance to their nearest known member and their burst rate at their typically observed energies. We extrapolate their rates using Eq. 1 to a common reference energy of $10^{29}\mbox{erg Hz}^{-1}$ from the closest available observed energy as well as to a central frequency of $\nu=600$MHz. In addition, we include rate estimates for FRBs 20181030A, 20121102A, 20201124A, 20220912A and 20241104A—sources not covered in Lu et al. (2022)—based on the findings of Li et al. (2021); Zhang et al. (2022, 2023); Chime/Frb Collaboration et al. (2023); Kirsten et al. (2024); Ould-Boukattine et al. (2024); Konijn et al. (2024); Tian et al. (2025); Shin et al. (2025). For SGR sources, we use updated rates from GReX observations (Shila et al., 2025). The resulting distribution is shown in Fig. 1. While individual rate estimates are subject to uncertainties - and, as discussed in §1, some sources show temporal variability and deviate from a single PL energy distribution - the aim of Fig. 1 is to capture the overall population trends using a simple model that avoids over-fitting. A striking feature of this plot is that the different FRB sub-classes span roughly nine orders of magnitude in number density and eight in burst rate (extrapolated to the fixed reference energy). Remarkably, the sources lie along a PL-like trend with $n_{\rm obs}\!\propto\!\mathcal{R}^{-1}$, in agreement with Eq. 20 and suggesting $a\approx 1$.
@@
Figure 1: Estimated source number densities of different FRB sub-classes. These are plotted against the rate per source extrapolated to an intermediate common energy of $10^{28}\mbox{erg Hz}^{-1}$ using the measured $\gamma$ or $\gamma\!\approx\!1.7\!\pm\!0.2$ when not available. A diamond indicates the same properties under the assumption (disfavored by our analysis) that a single source type dominates CHIME data. Since the latter is dominated by non-repeaters, this also provides a rough estimate for the number density and rate of CHIME non-repeaters. We find that $n_{\rm obs}\!\mathrel{\raisebox{3.01385pt}{\mbox{$\underset{\vbox{\hbox{$\sim$}}}{\vbox{\hbox{$\propto$}}}$}}}\!\mathcal{R}^{-1}$ over a factor of $\sim\!10^{9}$ in density, suggesting that the large range between common/inactive and rare/active sources might be fit by a single PL distribution with roughly equal contributions at different logarithmic scales.
@@
An alternative, related possibility, is that sources that are more active at a particular energy, may also be able to produce more energetic bursts, i.e. that $\mathcal{R}_{*}$ correlates with $E_{\rm c}$. Such a possibility, and physical motivation for it, is described in appendix §B. As it unavoidably introduces additional apriori free parameters, we apply the principle of Occam’s razor and focus primarily on the case with no such correlation. Future data will determine whether this complication is essential.
@@
For a variable rate, the results remain qualitatively similar so long as $k\!\lesssim\!a$, such that the most common sources, with $\mathcal{R}_{\rm*}\!\approx\!\mathcal{R}_{\rm 0,*}$ dominate. Since $E_{\rm all,k}$ increases with $R_{\rm*}$, for $k\!>\!a$ and $r\!>\!r_{z}\sqrt{\Phi_{\rm all,k}(\mathcal{R}_{\rm 0,*})/\Phi}$, sources with a rate such that $E_{\nu}\!=\!E_{\rm all,k}(\mathcal{R}_{\rm*})$ dominate. This leads to $\mathcal{R}_{\rm*}\!\propto\!r^{2(\gamma-1)}$ and using Eq. 3 to $\frac{\partial N_{\geq k}(>\Phi)}{\partial\log r}\propto r^{3-2a(\gamma-1)}$. The implication is that there is a critical value of $a\!\equiv\!a_{\gamma}\!=\!k_{\gamma}$, such that for $a\!<\!a_{\gamma}\&k$, most observed sources are distant and rare, while for $k\!>\!a\!>\!a_{\gamma}$, sources seen with $k$ times are preferentially nearby and are associated with intrinsically active objects. The results of the fluence and distance distributions with variable source rates, as well as the qualitative behavior in the $a-k$ parameter space, are illustrated in Fig. 2. They can be divided into three cases:
@@
    $a\!<\!a_{\gamma}\!=\!k_{\gamma}$ The least active of the observed sources (with $k\!<\!a$), are common (i.e. typical activity rate) and far ($r\!\sim\!r_{z}$). For larger $k\!>\!a$, the typical observed sources are instead dominated by rare (intrinsically highly active) and far sources. In this case $N_{\geq k}(>\!\Phi)\!\propto\!(\Phi/\Phi_{\rm all,k})^{a(1-\gamma)}\!\propto\!k^{-a}$ (or equivalently $N_{k}(>\!\Phi)\!\propto\!k^{-a-1}$). The number of sources with a given repetition number falls off slower than for the constant rate case.
@@
    $a\!>\!a_{\gamma}\!=\!k_{\gamma}$ Sources observed with $k\!<\!k_{\gamma}$ are common (i.e. typical activity rate) and far ($r\!\sim\!r_{z}$). Sources seen $k_{\gamma}\!<\!k\!<\!a$ are common and nearby ($r\!\ll\!r_{z}$). Finally, the sources seen to repeat most often, with $k\!>\!a\!>\!k_{\gamma}$, are both rare and nearby.
@@ 4 Inferences from observations
With approximately one MW-like galaxy per $100\mbox{Mpc}^{3}$, we can estimate the number density of FRB-producing magnetars in the local Universe as $10^{7}-3\times 10^{8}\mbox{Gpc}^{-3}$ (the range corresponding to either 1 or all magnetars in the Galaxy being potential FRB sources; for quantitative estimates below, we take the logarithmic mean). We use this to calculate the total number of FRB producing magnetars in the Universe, $N_{\rm src}\!\sim\!7\!\times\!10^{11}$, assuming that the magnetar density approximately tracks the star formation rate. At the same time, the number of sources seen by CHIME once or more during the period of data collection for the first CHIME catalog is only, $N_{\geq 1,\rm CH}=492$ and the number of sources observed at least twice (repeaters), $N_{\geq 2,\rm CH}=16$ (CHIME/FRB Collaboration et al., 2021). Correcting for the fraction of the sky viewable by CHIME, $f_{\rm FoV}\approx 0.5$, we see that
$\frac{N_{\rm\geq 1,CH}}{N_{\rm\geq 0,CH}}\approx 1.3\times 10^{-9\pm 0.74}\quad;\quad\frac{N_{\rm\geq 2,CH}}{N_{\rm\geq 1,CH}}\approx 0.03$ (22)
@@
Consider first the case of a single source type. In this case, since $N_{\geq 2}/N_{\geq 1}\ll 1$, then $\lambda_{\rm lim}\ll 1$, which using Eq. 14 leads to
@@
which is too large by at least seven orders of magnitude (see Eq. 22)  [^8] Accounting for the curvature of $N_{\geq k}(>\Phi)$ as well as for the burst rate as a function of frequency, cosmological corrections and a more realistic survey observation time distribution (§A), decreases $N_{\geq 1}/N_{\geq 0}$ by a factor of $\sim 26$ for a given $N_{\geq 2}/N_{\geq 1}$. This is still many orders of magnitude too large to accommodate SGR-type sources and typical CHIME sources as part of a single distribution..
@@
and $\lambda_{\rm lim}\approx\lambda_{\rm 0,lim}$ as before. For $a<1$, the fraction of repeaters is too large $N_{\geq 2}/N_{\geq 1}\geq 2^{-a}>1/2$, while for $a>2$, the estimate in Eq. 23 is approximately applicable, and once more the total number of sources is much too low. However, for intermediate values of $21<a<$ 2, and specifically, for $a-1\ll 1$, one finds that $N_{\geq 1}/N_{\geq 0}\approx\lambda_{\rm 0,lim}\ll N_{\geq 2,\rm CH}/N_{\geq 1,\rm CH}$. Indeed taking $a\approx 1.13$ simultaneously reproduces all three observables. As mentioned above, the simple analytic estimate needs to be adjusted to account for a few effects: (i) A more realistic, not broken PL approximation of $N_{\geq k}(>\Phi)$, (ii) cosmological distance, time dilation and frequency redshift corrections, (iii) The average spectrum of FRB sources, suggesting that bursts are less common at high frequencies and (iv) the non-uniform observation time spent by a large sky survey per point in the observed sky. These factors are all accounted for by direct numerical integration in §A. The results are shown in Fig. 3. As seen in the figure, the analytic result quoted above is well reproduced by this more precise calculation. The results of the latter, indicate that $1.1\lesssim a\lesssim 1.3$ can simultaneously match the observed $N_{\geq 1}/N_{\geq 0}$, $N_{\geq 2}/N_{\geq 1}$ and $\lambda_{\rm 0,lim}$.
@@
An alternative possible resolution is to assume a connection between source rate and maximum burst energy, the results corresponding to this situation are described in §B.
@@ 5 Implications for magnetar sources
Eq. 1 that has been applied here to the rate of bursts per source, has also been shown to hold across the Galactic magnetar population, and across temporally removed burst episodes in individual magnetars (see e.g. Cheng et al. 1996; Göǧüş et al. 1999; Gavriil et al. 2004). Moreover, these studies have found the PL index to be $\gamma\approx 1.7$, consistent with the values inferred for FRBs [^9] Magnetar bursts are thought to be associated with crust dislocations (Thompson & Duncan, 2001). If, with more observations, it is established that the value of $\gamma$ for FRBs is slightly different than that for magnetar bursts, this could indicate a difference in the scaling of X-ray/$\gamma$-ray bursts and that of FRBs with respect to the underlying distribution of dislocation amplitudes (see details in Wadiasingh et al. 2020)..
@@
For $\gamma<2$ the emitted power per source (averaged over sufficiently long timescales) is dominated by bursts with energies comparable to the maximal energy, $E_{\rm c}$. The (isotropic equivalent) radio energy released per source during an activity time $t$ is given by $E_{\rm rel}\approx 6\times 10^{36}(\mathcal{R}_{*}/\mathcal{R}_{0,*})(t/30\mbox{ yr})(E_{\rm c}/E_{\rm lim})^{2-\gamma}\mbox{ erg}$. Since we infer $|a-1|\ll 1$, there is large probability of detecting sources with $\mathcal{R}_{*}\gg\mathcal{R}_{0,*}$ and indeed such sources should dominate the population of CHIME repeaters (see Fig. 2). As shown in Fig. 1, for R1-type sources, $\mathcal{R}_{*}/\mathcal{R}_{0,*}\approx 3\times 10^{8}$. This leads to $E_{\rm rel,R1}\approx 1.8\times 10^{45}$ for the same beaming factor, radio efficiency and activity time, as consistent with the independent estimate $E_{\rm rel,R1}\approx 8.4\times 10^{44}\mbox{ erg}$ done in Beniamini & Kumar (2025)101010the small difference arises mostly due to uncertainty in the assumed spectral index and $\gamma$, which may also vary somewhat between R1 and the average FRB population. For an extremely active source such as R1, Beniamini & Kumar (2025) have shown that the magnetic energy reservoir becomes constraining and special conditions have to be met to satisfy it (e.g. regarding the alignment of magnetic, rotation and viewing angle axes as well as the age of the source and its internal magnetic field strength).

### v2

@@ 3 Varying activity rates
Crucially, the assumption of a PL rate distribution is directly supported by observations. Lu et al. (2022) estimated the number density of several empirically defined FRB sub-classes, based on the distance to their nearest known member and their burst rate at their typically observed energies. We extrapolate their rates using Eq. 1 (with $\gamma=1.7$ unless otherwise measured for the same source) to a common reference energy of $10^{28}\mbox{erg Hz}^{-1}$ from the closest available observed energy as well as to a central frequency of $\nu=600$MHz using $E_{\nu}\propto\nu^{\alpha}$ with $\alpha=-1.5$ (Macquart et al., 2019; Shin et al., 2023), unless otherwise measured from the same source. In addition, we include rate estimates for FRBs 20181030A, 20121102A, 20201124A, 20220912A, 20240114A, 20171020 and 20250316A - sources not covered in Lu et al. (2022) - based on the findings of Shannon et al. (2018); Li et al. (2021); Zhang et al. (2022, 2023); Chime/Frb Collaboration et al. (2023); Kirsten et al. (2024); Ould-Boukattine et al. (2024); Konijn et al. (2024); Tian et al. (2025); Shin et al. (2025); FRB Collaboration et al. (2025); Zhang et al. (2025). For SGR sources, we use updated rates from GReX observations (Shila et al., 2025). More details regarding the methodology of producing the figure are given in appendix B.
The resulting distribution is shown in Fig. 1. While individual rate estimates are subject to uncertainties - and, as discussed in §1, some sources show temporal variability and deviate from a single PL energy distribution - the aim of Fig. 1 is to capture the overall population trends using a simple model that avoids over-fitting. A striking feature of this plot is that the different FRB sub-classes span roughly nine orders of magnitude in number density and eight in burst rate (extrapolated to the fixed reference energy). Remarkably, the sources lie along a PL-like trend with $n_{\rm obs}\!\propto\!\mathcal{R}^{-1}$, in agreement with Eq. 20 and suggesting $a\approx 1$. As a comparison, we include in Fig.1, a point representing CHIME detected sources for the hypothetical test case (disfavored in this work) that there is a single activity rate per source underlying all sources (as per §2). Details of this calculation are given in appendix B.
The trend $n_{\rm obs}\propto\mathcal{R}^{-a}$ implies $\mathcal{R}\propto d_{\rm p}^{3/a}$ (where $d_{\rm p}$ is the proper distance). When extending the analysis to a larger group of sources, including those that are not the nearest known members of their population sub-class, one therefore expects known sources to span the ‘triangular’ range in the distance - activity rate ($d_{\rm p}$-$\mathcal{R}$) plane up to the $\mathcal{R}\propto d_{\rm p}^{3/a}$ trend line. Indeed, this too is reproduced, as shown in the bottom panel of Fig. 1. This supports the hypothesis of a single source population with a PL distribution of activity levels. That being said, we stress that the possibility of distinct multiple source populations cannot be ruled out - they are simply not required by these observational constraints.
@@
Figure 1: Top: Estimated source number densities of different FRB sub-classes. These are plotted against the rate per source extrapolated to an intermediate common energy of $10^{28}\mbox{erg Hz}^{-1}$ using the measured $\gamma$ or $\gamma\!\approx\!1.7\!\pm\!0.2$ when not available. Uncertainties mark the range corresponding to $90\%$ confidence limits. A diamond indicates the same properties for the hypothetical test case (disfavored by our analysis, see §B for details) that a single source type dominates CHIME data. Since the latter is dominated by non-repeaters, this also provides a rough estimate for the number density and rate of CHIME non-repeaters. We find that approximately $n_{\rm obs}\!\propto\!\mathcal{R}^{-1}$ over a factor of $\sim\!10^{9}$ in density, suggesting that the large range between common/inactive and rare/active sources might be fit by a single PL distribution with roughly equal contributions at different logarithmic scales. Bottom: Rate per source as a function of (proper) distance. Results are shown for extragalactic FRB repeaters and non-repeaters with well constrained distances and rates. Downward facing triangles indicate $90\%$ upper limits. Consistency with the PL distribution of activity rates, $n_{\rm obs}\propto\mathcal{R}^{-a}$, requires that sources occupy the region below a boundary given by $\mathcal{R}\propto d_{\rm p}^{3/a}$ - i.e. within the blue shaded region.
@@
An alternative, related possibility, is that sources that are more active at a particular energy, may also be able to produce more energetic bursts, i.e. that $\mathcal{R}_{*}$ correlates with $E_{\rm c}$. Such a possibility, and physical motivation for it, is described in appendix §C. As it unavoidably introduces additional apriori free parameters, we apply the principle of Occam’s razor and focus primarily on the case with no such correlation. Future data will determine whether this complication is essential.
@@
For a variable rate, the results remain qualitatively similar so long as $k\!\lesssim\!a$, such that the most common sources, with $\mathcal{R}_{\rm*}\!\approx\!\mathcal{R}_{\rm 0,*}$ dominate. Since $E_{\rm all,k}$ increases with $\mathcal{R}_{\rm*}$, for k>ak\!>\!a and $r\!>\!r_{z}\sqrt{\Phi_{\rm all,k}(\mathcal{R}_{\rm 0,*})/\Phi}$, sources with a rate such that $E_{\nu}\!=\!E_{\rm all,k}(\mathcal{R}_{\rm*})$ dominate. This leads to $\mathcal{R}_{\rm*}\!\propto\!r^{2(\gamma-1)}$ and using Eq. 3 to $\frac{\partial N_{\geq k}(>\Phi)}{\partial\log r}\propto r^{3-2a(\gamma-1)}$. The implication is that there is a critical value of $a\!\equiv\!a_{\gamma}\!=\!k_{\gamma}$, such that for $a\!<\!a_{\gamma}$ and $a<k$, most observed sources are distant and rare, while for $k\!>\!a\!>\!a_{\gamma}$, sources seen $k$ times are preferentially nearby and are associated with intrinsically active objects. The results of the fluence and distance distributions with variable source rates, as well as the qualitative behavior in the aka-k parameter space, are illustrated in Fig. 2. They can be divided into three cases:
@@
    $a\!<\!a_{\gamma}\!=\!k_{\gamma}$ - The least active of the observed sources (with k<ak\!<\!a), are common (i.e. typical activity rate) and far ($r\!\sim\!r_{z}$). For larger k>ak\!>\!a, the typical observed sources are instead dominated by rare (intrinsically highly active) and far sources. In this case $N_{\geq k}(>\!\Phi)\!\propto\!(\Phi/\Phi_{\rm all,k})^{a(1-\gamma)}\!\propto\!k^{-a}$ (or equivalently $N_{k}(>\!\Phi)\!\propto\!k^{-a-1}$). The number of sources with a given repetition number falls off slower than for the constant rate case.
@@
    $a\!>\!a_{\gamma}\!=\!k_{\gamma}$ - Sources observed with $k\!<\!k_{\gamma}$ are common (i.e. typical activity rate) and far ($r\!\sim\!r_{z}$). Sources seen $k_{\gamma}\!<\!k\!<\!a$ are common and nearby ($r\!\ll\!r_{z}$). Finally, the sources seen to repeat most often, with $k\!>\!a\!>\!k_{\gamma}$, are both rare and nearby.
@@ 4 Inferences from observations
With approximately one Milky Way-like galaxy per $100\mbox{Mpc}^{3}$(Leroy et al., 2019) and 30 known active magnetars in the Milky Way (Kaspi & Beloborodov, 2017), we can estimate the number density of FRB-producing magnetars in the local Universe as $10^{7}-3\times 10^{8}\mbox{Gpc}^{-3}$ (the range corresponding to either 1 or all magnetars in the Galaxy being potential FRB sources; for quantitative estimates below, we take the logarithmic mean). We use this to calculate the total number of FRB producing magnetars in the Universe, $N_{\rm src}\!\sim\!7\times\!10^{11}$, assuming that the magnetar density approximately tracks the star formation rate. At the same time, the number of sources seen by CHIME once or more during the period of data collection for the first CHIME catalog is only, $N_{\geq 1,\rm CH}^{\rm obs}=492$ and the number of sources observed at least twice (repeaters), $N_{\geq 2,\rm CH}^{\rm obs}=16$ (CHIME/FRB Collaboration et al., 2021). Correcting for the fraction of the sky viewable by CHIME, $f_{\rm srv}\approx 0.5$ (i.e. the all-sky equivalent $N_{\geq 1,\rm CH}^{\rm all-sky}$ is given from the observed number $N_{\geq 1,\rm CH}^{\rm obs}$ via $N_{\geq 1,\rm CH}^{\rm all-sky}$$N_{\geq 1,\rm CH}^{\rm obs}/f_{\rm srv}$), we see that
$\frac{N_{\rm\geq 1,CH}^{\rm all-sky}}{N_{\rm\geq 0,CH}^{\rm all-sky}}\approx 1.3\times 10^{-9\pm 0.74}\quad;\quad\frac{N_{\rm\geq 2,CH}^{\rm all-sky}}{N_{\rm\geq 1,CH}^{\rm all-sky}}\approx 0.03$ (22)
@@
Consider first the case of a single source type. Since $N_{\geq 2}/N_{\geq 1}\!\ll\!1$, then $\lambda_{\rm lim}\!\ll\!1$, which using Eq. 14 leads to
@@
which is too large by at least seven orders of magnitude (see Eq. 22)  [^9] Accounting for the curvature of $N_{\geq k}(>\Phi)$ as well as for the burst rate as a function of frequency, cosmological corrections, a more realistic survey observation time distribution (§A) and a non-uniform density of sources, decreases $N_{\geq 1}/N_{\geq 0}$ by a factor of $\sim 26$ for a given $N_{\geq 2}/N_{\geq 1}$. This is still many orders of magnitude too large to accommodate SGR-type sources and typical CHIME sources as part of a single distribution..
@@
and $\lambda_{\rm lim}\approx\lambda_{\rm 0,lim}$ as before. For $a<1$, the fraction of repeaters is too large $N_{\geq 2}/N_{\geq 1}\geq 2^{-a}>1/2$, while for $a>2$, the estimate in Eq. 23 is approximately applicable, and once more the total number of sources is much too low. However, for intermediate values of $1<a<2$, and specifically, for $a-1\ll 1$, one finds that $N_{\geq 1}/N_{\geq 0}\approx\lambda_{\rm 0,lim}\ll N_{\geq 2,\rm CH}^{\rm all-sky}/N_{\geq 1,\rm CH}^{\rm all-sky}$. Indeed taking $a\approx 1.13$ simultaneously reproduces all three observables. As mentioned above, the simple analytic estimate needs to be adjusted to account for a few effects: (i) A more realistic, not broken PL approximation of $N_{\geq k}(>\Phi)$, (ii) cosmological distance, time dilation and frequency redshift corrections, (iii) The average spectrum of FRB sources, suggesting that bursts with a given spectral flux are less common at high frequencies, (iv) the non-uniform observation time spent by a large sky survey per point in the observed sky and (v) the number density of sources which in general depends on redshift. These factors are all accounted for by direct numerical integration in §A. The results are shown in Fig. 3. As seen in the figure, the analytic result quoted above is well reproduced by this more precise calculation. The results of the latter, indicate that $1.1\lesssim a\lesssim 1.3$ can simultaneously match the observed $N_{\geq 1}/N_{\geq 0}$, $N_{\geq 2}/N_{\geq 1}$ and $\lambda_{\rm 0,lim}$.
@@
An alternative possible resolution is to assume a connection between source rate and maximum burst energy, the results corresponding to this situation are described in §C.
@@ 5 Implications for magnetar sources
Our analysis outlined in §4 and appendix A, is model independent and does not rely on FRBs being associated with a specific source type or emission mechanism. Nonetheless, various lines of evidence point towards magnetars being the underlying sources of a large fraction, and perhaps all of FRBs. In particular, Eq. 1 that has been applied here to the rate of FRBs per source, has also been shown to hold across the Galactic magnetar population, and across temporally removed burst episodes in individual magnetars (see e.g. Cheng et al. 1996; Göǧüş et al. 1999; Gavriil et al. 2004). Moreover, these studies have found the PL index to be $\gamma\approx 1.7$, consistent with the values inferred for FRBs [^10] Magnetar bursts are thought to be associated with crust dislocations (Thompson & Duncan, 2001). If, with more observations, it is established that the value of $\gamma$ for FRBs is slightly different than that for magnetar bursts, this could indicate a difference in the scaling of X-$\gamma-ray$ bursts and that of FRBs with respect to the underlying distribution of dislocation amplitudes (see details in Wadiasingh et al. 2020).. Other lines of evidence include radiation properties (coherent emission, rapid variability, narrow spectral width, bright luminosity, see Popov & Postnov 2010; Kumar et al. 2017; Beniamini & Kumar 2020) as well as FRB host galaxies, offsets relative to their hosts and burst statistics (Wadiasingh & Timokhin, 2019; Cheng et al., 2020; Gordon et al., 2023; Totani & Tsuzuki, 2023). The most direct association comes from the discovery of FRB 20200428, originating from a well-known Galactic magnetar (CHIME/FRB Collaboration et al., 2020; Bochenek et al., 2020). However, this is also a unique source, in that its rate (at a given energy) is very low and the number density of similar sources in the Universe is very large (hence its location in the top left of figure 1). This then raises the question of whether magnetars can potentially accommodate the entire range of FRB source activity levels.
As described in §4, when calibrating our framework by the FRB rate per source and volumetric density of Galactic magnetars, we are able to simultaneously reconcile the ratio of observed non-repeaters compared to all SGRs, the ratio of observed repeaters to non-repeaters and the typical distances of detected repeaters, being closer by a few tens of percent than those of detected non-repeaters. This is done for a population with a continuous PL function describing the probability of the source having a given activity level ($P(\geq\mathcal{R}_{*})\propto\mathcal{R}_{*}^{-a}$, see Eq. 20). Remarkably, once calibrating the rate and density of the most common sources as described above, the only free parameter of the model is $a$, and the range found to reproduce the observables is obtained for the same value of $a$ that is independently obtained from observations of different individual sources (see figure 1 and appendix B). This argues in favor of magnetars being the source of a large fraction (or all) of the FRB population.
@@
For $\gamma<2$, inferred for FRB sources and magnetar X-ray / $\gamma$-ray bursts, the emitted power per source (averaged over sufficiently long timescales) is dominated by bursts with energies comparable to the maximal energy, $E_{\rm c}$. The (isotropic equivalent) radio energy released per source during an activity time $t$ is given by $E_{\rm rel}\approx 6\times 10^{36}(\mathcal{R}_{*}/\mathcal{R}_{0,*})(t/30\mbox{ yr})(E_{\rm c}/E_{\rm lim})^{2-\gamma}\mbox{ erg}$. Since we infer $|a-1|\ll 1$, there is large probability of detecting sources with $\mathcal{R}_{*}\gg\mathcal{R}_{0,*}$ (despite their intrinsic rarity) and indeed such sources should dominate the population of CHIME repeaters (see Fig. 2). As shown in Fig. 1, for FRB 20121102A-type sources, $\mathcal{R}_{*}/\mathcal{R}_{0,*}\approx 3\times 10^{8}$. This leads to $E_{\rm rel}\approx 1.8\times 10^{45}$ for the same beaming factor, radio efficiency and activity time, as consistent with the independent estimate $E_{\rm rel}\approx 8.4\times 10^{44}\mbox{ erg}$ done in Beniamini & Kumar (2025)111111the small difference arises mostly due to uncertainty in the assumed spectral index and $\gamma$, which may also vary somewhat between FRB 20121102A and the average FRB population. For an extremely active source such as FRB 20121102A, Beniamini & Kumar (2025) have shown that the magnetic energy reservoir becomes constraining and special conditions have to be met to satisfy it (e.g. regarding the alignment of magnetic, rotation and viewing angle axes as well as the age of the source and its internal magnetic field strength). The apparent source density of FRB 20121102A-like sources is extremely small, $\sim 10^{-9}$ that of standard SGRs (see figure 1). This can be a combination of intrinsic magnetar properties (e.g. extremely strong and rare magnetic field strength or very young age) as well as strong beaming (e.g. due to alignment of the magnetic ans spin axes, see Beniamini & Kumar 2025). We return in more detail to the potential role of the latter in §6.